OpenStudy (anonymous):

Which of the following is a quadratic function? Answer y = x3 − 3x − 20 y = 3x – x + 4 y = x2 – 4x + 7 y = 2x4 + 9x + 17

OpenStudy (anonymous):

y = x2-4x+7

OpenStudy (lgbasallote):

quadratic function means the degree is 2

OpenStudy (lgbasallote):

which of the choices has degree 2?

OpenStudy (anonymous):

Note that a quadratic function has the terms: a(x^2) + bx + c - one with x^1 (or x) and (x^2). Therefore, C is the correct answer.

OpenStudy (lgbasallote):

not necessarily it has one x^1

OpenStudy (lgbasallote):

\[y = x^2+1\] that's still quadratic \[y = x^2\] that's also quadratic \[y = x+1\] that however is not quadratic anymore